OpenStudy (anonymous):

If f(x) = 2x and g(x) = x-4, what is the value of f[g(3)]?

OpenStudy (anonymous):

What you wanna do is plug 3 into g(x)=x-4, then plug that value into F(x)=2x

OpenStudy (anonymous):

g(3)=3-4 =-1, Then f(-1)=2(-1) =-2